The ratio of two numbers is 5 to 3, and their sum is 56. What are the numbers?
Answer by stanbon(75887)

The ratio of two numbers is 5 to 3, and their sum is 56. What are the numbers?
---
5 to 3 is the same as 5x to 3x
-----
Equation:
5x + 3x = 56
7x = 56
x = 8
-------
5x = 5*8 = 40
3x = 3*8 = 24
=================
Cheers,
Stan H.
=================